OpenStudy (anonymous):

Help! I'll give you a medal! Percy started a new job and must enroll in a new family health insurance plan. One of the options involves prescription drug coverage. Percy estimates that he and his family will fill 4 prescriptions per month, totaling $450. Which option would result in the lowest overall cost for Percy? $120 monthly premium; $14 copayment per prescription $160 monthly premium; $10 copayment per prescription $50 monthly premium; 80% coverage for first $500 in prescription costs, then 90% coverage for all prescription costs over $500 $86 monthly premium; 70% cov

OpenStudy (anonymous):

$86 monthly premium; 70% coverage for all prescription costs

OpenStudy (anonymous):

For choice A , I got $175.00 , Choice B $200 Choice D $221.00 Choice C I need help on . I don't know how to solve for it! Help please

OpenStudy (anonymous):

50 + (.80*450) = 410

OpenStudy (anonymous):

a. 120 + (14*4) = 120 + 56 = 176 b. 160+ (10*4) = 160 + 40 = 200 c . 50 + (.80*450) = 410 d. 86 + .70*450 = 86 +315 = 401 note these are all monthly costs... so you should multiply by 12 to determine the annual cost